OpenStudy (kaley246):

Which of the following is the vertex of the function? y = x2 - 6x + 9 (Points : 4) (3, 0) (-3, 36) (4, 1) (1, 4)

OpenStudy (anonymous):

use the discriminant to find the x value of the vertex x = -b/2a [y = ax^2 + bx + c] x = -(-6)/2(1) = 6/2 = 3 now that we know the x coordinate is 3 we can plug that back into the original equation to find the y coordinate. y = (3)^2 - 6(3) + 9 = 9 - 18 + 9 = 0 so the vertex is at (3,0)